OpenStudy (anonymous):

(0, -2) is a solution of the equation 2x - 3y = 6 True or False?

OpenStudy (anonymous):

True

OpenStudy (muzzack):

replace the equation \[\Large\cal\ 2(0) - 3(-2) = 6\]

OpenStudy (muzzack):

True

OpenStudy (anonymous):

(0,-2) is an answer to the equation \[2x-3y=6\] \[2(0)-3(-2)=6\] \[6=6\]